Let f(x)=-2x-7 and g(x)=-4x+3. Find (fog)(-5)
Answer by sudhanshu_kmr(1152)


g(-5) = -4 * -5 + 3 = 20 +3 = 23
f(g(5))
= f(23) = -2 * 23 -7
= -46 -7
= -53